# Demonstration of Hardware Efficient Photonic Variational Quantum Algorithm

### v1

@@ Acknowledgements
This research was funded in whole, or in part, by Horizon 2020 and Horizon Europe research and innovation programme under grant agreement No 899368 (EPIQUS) and No 101135288 (EPIQUE), the Marie Skłodowska-Curie grant agreement No 956071 (AppQInfo), and the QuantERA II Programme under Grant Agreement No 101017733 (PhoMemtor). Further funding was received from the Austrian Science Fund (FWF) through Quantum Science Austria (COE1), BeyondC (Grant-DOI 10.55776/F71) and Research Group 5(FG5); and from the Austrian Federal Ministry for Digital and Economic Affairs, the National Foundation for Research, Technology and Development and the Christian Doppler Research Association. K.P. and X.C. also appreciate the funding from the Basque Government through Grant No. IT1470-22, the project grant PID2021-126273NB-I00 funded by MCIN/AEI/10.13039/501100011033 and by “ERDF A way of making Europe” and “ERDF Invest in your Future”, and the Spanish Ministry of Economic Affairs and Digital Transformation through the QUANTUM ENIA project call-Quantum Spain project. For the purpose of open access, the author has applied a CC BY public copyright licence to any Author Accepted Manuscript version arising from this submission.

### v2

@@ Acknowledgements
This research was co-funded by the European Union (HORIZON Europe Research and Innovation Programme, EPIQUE, No 101135288). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Commission-EU. Neither the European Union nor the granting authority can be held responsible for them. This project has received funding from the European Union’s Horizon 2020 research and innovation programme under grant agreement No 899368 (EPIQUS), under the Marie Skłodowska-Curie grant agreement No 956071 (AppQInfo) and under Grant Agreement no. 101017733 (QuantERA II). This research was funded in whole or in part by the Austrian Science Fund (FWF)[10.55776/ESP205] (PREQUrSOR), [10.55776/I6002] and [10.55776/F71] (BeyondC). For open access purposes, the author has applied a CC BY public copyright license to any author accepted manuscript version arising from this submission. The financial support by the Austrian Federal Ministry of Labour and Economy, the National Foundation for Research, Technology and Development and the Christian Doppler Research Association is gratefully acknowledged. K.P. and X.C. also appreciate the funding from the Basque Government through Grant No. IT1470-22, the project grant PID2021-126273NB-I00 funded by MCIN/AEI/10.13039/501100011033 and by “ERDF A way of making Europe” and “ERDF Invest in your Future”, and the Spanish Ministry of Economic Affairs and Digital Transformation through the QUANTUM ENIA project call-Quantum Spain project.